BRICS Research Series, 2001

August 5, 2003

This document is also available as PostScript and DVI.

References

[RS-01-55] Abstract, PostScript, PDF, DVI.

Daniel Damian and Olivier Danvy. A Simple CPS Transformation of Control-Flow Information. December 2001. 18 pp. Appears in Logic Journal of the IGPL, 10(2):501–515, 2002.

[RS-01-54] Abstract, PostScript, PDF, DVI.

Daniel Damian and Olivier Danvy. Syntactic Accidents in Program Analysis: On the Impact of the CPS Transformation. December 2001. 41 pp. To appear in the Journal of Functional Programming. This report supersedes the earlier BRICS report RS-00-15.

[RS-01-53] Abstract, PostScript, PDF, DVI.

Zoltn sik and Masami Ito. Temporal Logic with Cyclic Counting and the Degree of Aperiodicity of Finite Automata. December 2001. 31 pp.

[RS-01-52] Abstract, PostScript, PDF, DVI.

Jens Groth. Extracting Witnesses from Proofs of Knowledge in the Random Oracle Model. December 2001. 23 pp.

[RS-01-51] Abstract, PostScript, PDF, DVI.

Ulrich Kohlenbach. On Weak Markov’s Principle. December 2001. 10 pp. Appears in Mathematical Logic Quarterly, 48(suppl. 1):59–65, 2002.

[RS-01-50] Abstract, PostScript, PDF, DVI.

Jiř Srba. Note on the Tableau Technique for Commutative Transition Systems. December 2001. 19 pp. Appears in Nielsen and Engberg, editors, Foundations of Software Science and Computation Structures, FoSSaCS ’02 Proceedings, LNCS 2303, 2002, pages 387–401.

[RS-01-49] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Lasse R. Nielsen. A First-Order One-Pass CPS Transformation. December 2001. 21 pp. To appear in Theoretical Computer Science. Extended version of a paper appearing in Nielsen and Engberg, editors, Foundations of Software Science and Computation Structures, FoSSaCS ’02 Proceedings, LNCS 2303, 2002, pages 98–113.

[RS-01-48] Abstract, PostScript, PDF, DVI. Mogens Nielsen and Frank D. Valencia. Temporal Concurrent Constraint Programming: Applications and Behavior. December 2001. 36 pp. Appears in Brauer, Ehrig, Karhumki and Salomaa, editors, Formal and Natural Computing, LNCS 2300, 2001, pages 298–321.

[RS-01-47] Abstract, PostScript, PDF, DVI. Jesper Buus Nielsen. Non-Committing Encryption is Too Easy in the Random Oracle Model. December 2001. 20 pp.

[RS-01-46] Abstract, PostScript, PDF, DVI. Lars Kristiansen. The Implicit Computational Complexity of Imperative Programming Languages. November 2001. 46 pp.

[RS-01-45] Abstract, PostScript, PDF. Ivan B. Damgrd and Gudmund Skovbjerg Frandsen. An Extended Quadratic Frobenius Primality Test with Average Case Error Estimates. November 2001. 43 pp.

[RS-01-44] Abstract, PostScript, PDF. M. Oliver Mller, Harald Rue, and Maria Sorea. Predicate Abstraction for Dense Real-Time Systems. November 2001. 27 pp. Appears in Asarin, Maler and Yovine, editors, Workshop on Theory and Practice of Timed Systems, TPTS ’02 Proceedings, ENTCS 65(6), 2002.

[RS-01-43] Abstract, PostScript, PDF, DVI. Ivan B. Damgrd and Jesper Buus Nielsen. From Known-Plaintext Security to Chosen-Plaintext Security. November 2001. 18 pp.

[RS-01-42] Abstract, PostScript, PDF, DVI. Zoltn sik and Werner Kuich. Rationally Additive Semirings. November 2001. 11 pp. Appears in Journal of Universal Computer Science, 8(2):173–183, 2002.

[RS-01-41] Abstract, PostScript, PDF, DVI. Ivan B. Damgrd and Jesper Buus Nielsen. Perfect Hiding and Perfect Binding Universally Composable Commitment Schemes with Constant Expansion Factor. October 2001. 43 pp. Appears in Yung, editor, Advances in Cryptology: 22nd Annual International Cryptology Conference, CRYPTO ’02 Proceedings, LNCS 2442, 2002, pages 581–596.

[RS-01-40] Abstract, PostScript, PDF, DVI. Daniel Damian and Olivier Danvy. CPS Transformation of Flow Information, Part II: Administrative Reductions. October 2001. 9 pp. To appear in the Journal of Functional Programming. Superseeded by the BRICS report RS-02-36.

[RS-01-39] Abstract, PostScript, PDF, DVI. Olivier Danvy and Mayer Goldberg. There and Back Again. October 2001. 14 pp. To appear in the proceedings of the 2002 ACM SIGPLAN International Conference on Functional Programming.

[RS-01-38] Abstract, PostScript, PDF, DVI. Zoltn sik. Free De Morgan Bisemigroups and Bisemilattices. October 2001. 13 pp. To appear in the journal Algebra Colloquium.

[RS-01-37] Abstract, PostScript, PDF, DVI. Ronald Cramer and Victor Shoup. Universal Hash Proofs and a Paradigm for Adaptive Chosen Ciphertext Secure Public-Key Encryption. October 2001. 34 pp.

[RS-01-36] Abstract, PostScript, PDF. Gerth Stlting Brodal, Rolf Fagerberg, and Riko Jacob. Cache Oblivious Search Trees via Binary Trees of Small Height. October 2001. 20 pp. Appears in Eppstein, editor, The Thirteenths Annual ACM-SIAM Symposium on Discrete Algorithms, SODA ’02 Proceedings, 2002, pages 39–48.

[RS-01-35] Abstract, PostScript, PDF, DVI. Mayer Goldberg. A General Schema for Constructing One-Point Bases in the Lambda Calculus. September 2001. 8 pp.

[RS-01-34] Abstract, PostScript, PDF. Flemming Friche Rodler and Rasmus Pagh. Fast Random Access to Wavelet Compressed Volumetric Data Using Hashing. August 2001. 31 pp. To appear in ACM Transactions on Graphics.

[RS-01-33] Abstract, PostScript, PDF. Rasmus Pagh and Flemming Friche Rodler. Lossy Dictionaries. August 2001. 14 pp. Short version appears in Meyer auf der Heide, editor, 9th Annual European Symposium on Algorithms, ESA ’01 Proceedings, LNCS 2161, 2001, pages 300–311.

[RS-01-32] Abstract, PostScript, PDF. Rasmus Pagh and Flemming Friche Rodler. Cuckoo Hashing. August 2001. 21 pp. Short version appears in Meyer auf der Heide, editor, 9th Annual European Symposium on Algorithms, ESA ’01 Proceedings, LNCS 2161, 2001, pages 121–133.

[RS-01-31] Abstract, PostScript, PDF, DVI. Olivier Danvy and Lasse R. Nielsen. Syntactic Theories in Practice. July 2001. 37 pp. Extended version of an article to appear in the informal proceedings of the Second International Workshop on Rule-Based Programming, RULE 2001 (Firenze, Italy, September 4, 2001). Superseeded by the BRICS report RS-02-4.

[RS-01-30] Abstract, PostScript, PDF, DVI. Lasse R. Nielsen. A Selective CPS Transformation. July 2001. 24 pp. Appears in Brookes and Mislove, editors, 27th Annual Conference on the Mathematical Foundations of Programming Semantics, MFPS ’01 Proceedings, ENTCS 45, 2001. A preliminary version appeared in Brookes and Mislove, editors, 17th Annual Conference on Mathematical Foundations of Programming Semantics, MFPS ’01 Preliminary Proceedings, BRICS Notes Series NS-01-2, 2001, pages 201–222.

[RS-01-29] Abstract, PostScript, PDF, DVI. Olivier Danvy, Bernd Grobauer, and Morten Rhiger. A Unifying Approach to Goal-Directed Evaluation. July 2001. 23 pp. Appears in New Generation Computing, 20(1):53–73, November 2001. A preliminary version appeared in Taha, editor, 2nd International Workshop on Semantics, Applications, and Implementation of Program Generation, SAIG ’01 Proceedings, LNCS 2196, 2001, pages 108–125.

[RS-01-28] Abstract, PostScript, PDF, DVI. Luca Aceto, Zoltn sik, and Anna Inglfsdttir. A Fully Equational Proof of Parikh’s Theorem. June 2001. 28 pp. Appears in RAIRO, Theoretical Informatics and Applications, 36(2):129–153, 2002, special issue devoted to selected papers from FICS 2001 (A. Labella ed.).

[RS-01-27] Abstract, PostScript, PDF, DVI. Mario Jos Cccamo and Glynn Winskel. A Higher-Order Calculus for Categories. June 2001. 24 pp. Appears in Boulton and Jackson, editors, Theorem Proving in Higher Order Logics: 14th International Conference, TPHOLs ’01 Proceedings, LNCS 2152, 2001, pages 136– 153.

[RS-01-26] Abstract, PostScript, PDF, DVI. Ulrik Frendrup and Jesper Nyholm Jensen. A Complete Axiomatization of Simulation for Regular CCS Expressions. June 2001. 18 pp.

[RS-01-25] Abstract, PostScript, PDF, DVI. Bernd Grobauer. Cost Recurrences for DML Programs. June 2001. 51 pp. Extended version of a paper appearing in Leroy, editor, Proceedings of the 6th ACM SIGPLAN International Conference on Functional Programming, 2001, pages 253–264.

[RS-01-24] Abstract, PostScript, PDF, DVI. Zoltn sik and Zoltn L. Nmeth. Automata on Series-Parallel Biposets. June 2001. 15 pp. Appears in Kuich, Rozenberg and Salomaa, editors, 5th International Conference, Developments in Language Theory, DLT ’01 Revised Papers, LNCS 2295, 2002, pages 217–227. Superseded by the later BRICS report RS-02-44.

[RS-01-23] Abstract, PostScript, PDF, DVI. Olivier Danvy and Lasse R. Nielsen. Defunctionalization at Work. June 2001. 45 pp. Extended version of an article appearing in Sndergaard, editor, 3rd International Conference on Principles and Practice of Declarative Programming, PPDP ’01 Proceedings, 2001, pages 162–174.

[RS-01-22] Abstract, PostScript, PDF, DVI. Zoltn sik. The Equational Theory of Fixed Points with Applications to Generalized Language Theory. June 2001. 21 pp. Appears in Kuich, Rozenberg and Salomaa, editors, 5th International Conference, Developments in Language Theory, DLT ’01 Revised Papers, LNCS 2295, 2002,pages 21–36.

[RS-01-21] Abstract, PostScript, PDF, DVI. Luca Aceto, Zoltn sik, and Anna Inglfsdttir. Equational Theories of Tropical Semirings. June 2001. 52 pp. To appear in Theoretical Computer Science. Extended abstracts of parts of this paper have appeared in Honsell and Miculan, editors, Foundations of Software Science and Computation Structures, FoSSaCS ’01 Proceedings, LNCS 2030, 2001, pages 42–56 and in Gaubert and Loiseau, editors, Workshop on Max-Plus Algebras and Their Applications to Discrete-Event Systems, Theoretical Computer Science, and Optimization, MAX-PLUS ’01 Proceedings, IFAC (International Federation of Automatic Control) IFAC Publications, 2001.

[RS-01-20] Abstract, PostScript, PDF, DVI. Catuscia Palamidessi and Frank D. Valencia. A Temporal Concurrent Constraint Programming Calculus. June 2001. 31 pp. Appears in Walsh, editor, 7th International Conference on Principles and Practice of Constraint Programming, CP ’01 Proceedings, LNCS 2239, 2001, pages 302–316.

[RS-01-19] Abstract, PostScript, PDF, DVI. Jiř Srba. On the Power of Labels in Transition Systems. June 2001. 23 pp. Full and extended version of Larsen and Nielsen, editors, Concurrency Theory: 12th International Conference, CONCUR ’01 Proceedings, LNCS 2154, 2001, pages 277–291.

[RS-01-18] Abstract, PostScript, PDF. Katalin M. Hangos, Zsolt Tuza, and Anders Yeo. Some Complexity Problems on Single Input Double Output Controllers. May 2001. 27 pp.

[RS-01-17] Abstract, PostScript, PDF. Claus Brabrand, Anders Mller, Steffan Olesen, and Michael I. Schwartzbach. Language-Based Caching of Dynamically Generated HTML. May 2001. 18 pp. Appears in World Wide Web Journal, 5(4):305–323, 2002.

[RS-01-16] Abstract, PostScript, PDF, DVI. Olivier Danvy, Morten Rhiger, and Kristoffer H. Rose. Normalization by Evaluation with Typed Abstract Syntax. May 2001. 9 pp. Appears in Journal of Functional Programming, 11(6):673–68, November 2000.

[RS-01-15] Abstract, PostScript, PDF, DVI. Luigi Santocanale. A Calculus of Circular Proofs and its Categorical Semantics. May 2001. 30 pp. Appears in Nielsen and Engberg, editors, Foundations of Software Science and Computation Structures, FoSSaCS ’02 Proceedings, LNCS 2303, 2002, pages 129–143.

[RS-01-14] Abstract, PostScript, PDF, DVI. Ulrich Kohlenbach and Paulo B. Oliva. Effective Bounds on Strong Unicity in L_{1}-Approximation. May 2001. 38 pp. A revised and numerically much improved version appeared in Annals of Pure and Applied Logic, 121:1–38, 2003. Extended extended abstract appears in Hofmann, editor, 3rd International Workshop on Implicit Computational Complexity, ICC ’01 Proceedings, BRICS Notes Series NS-01-3, 2001, pages 117–122.

[RS-01-13] Abstract, PostScript, PDF, DVI. Federico Crazzolara and Glynn Winskel. Events in Security Protocols. April 2001. 30 pp. Appears in 8th ACM conference on Computer and Communications Security, CCS ’01 Proceedings, 2001, pages 96–105.

[RS-01-12] Abstract, PostScript, PDF, DVI. Torben Amtoft, Charles Consel, Olivier Danvy, and Karoline Malmkjr. The Abstraction and Instantiation of String-Matching Programs. April 2001. 37 pp. Appears in Mogensen, Schmidt and Sudborough, editors, The Essence of Computation: Complexity, Analysis, Transformation, LNCS 2556, 2002, pages 332–357.

[RS-01-11] Abstract, PostScript, PDF. Alexandre David and M. Oliver Mller. From HUppaal to Uppaal: A Translation from Hierarchical Timed Automata to Flat Timed Automata. March 2001. 40 pp. Appears in Kutsche and Weber, editors, Fundamental Approaches to Software Engineering: Fifth International Conference, FASE ’02 Proceedings, LNCS 2306, 2002, pages 218–232, with the title Formal Verification of UML Statecharts with Real-Time Extensions.

[RS-01-10] Abstract, PostScript, PDF, DVI. Daniel Fridlender and Mia Indrika. Do we Need Dependent Types? March 2001. 6 pp. Appears in Journal of Functional Programming, 10(4):409–415, 2000. Superseeds BRICS Report RS-98-38.

[RS-01-9] Abstract, PostScript, PDF. Claus Brabrand, Anders Mller, and Michael I. Schwartzbach. Static Validation of Dynamically Generated HTML. February 2001. 18 pp. Appears in Field and Snelting, editors, ACM SIGPLAN-SIGSOFT Workshop on Program Analysis For Software Tools and Engineering, PASTE ’01 Proceedings, 2001, pages 38–45.

[RS-01-8] Abstract, PostScript, PDF. Ulrik Frendrup and Jesper Nyholm Jensen. Checking for Open Bisimilarity in the π-Calculus. February 2001. 61 pp.

[RS-01-7] Abstract, PostScript, PDF, DVI. Gregory Gutin, Khee Meng Koh, Eng Guan Tay, and Anders Yeo. On the Number of Quasi-Kernels in Digraphs. January 2001. 11 pp.

[RS-01-6] Abstract, PostScript, PDF, DVI. Gregory Gutin, Anders Yeo, and Alexey Zverovich. Traveling Salesman Should not be Greedy: Domination Analysis of Greedy-Type Heuristics for the TSP. January 2001. 7 pp.

[RS-01-5] Abstract, PostScript, PDF. Thomas S. Hune, Judi Romijn, Marille Stoelinga, and Frits W. Vaandrager. Linear Parametric Model Checking of Timed Automata. January 2001. 44 pp. Appears in Margaria and Yi, editors, Tools and Algorithms for The Construction and Analysis of Systems: 7th International Conference, TACAS ’01 Proceedings, LNCS 2031, 2001, pages 174–188.

[RS-01-4] Abstract, PostScript, PDF. Gerd Behrmann, Ansgar Fehnker, Thomas S. Hune, Kim G. Larsen, Paul Pettersson, and Judi Romijn. Efficient Guiding Towards Cost-Optimality in Uppaal. January 2001. 21 pp. Appears in Margaria and Yi, editors, Tools and Algorithms for The Construction and Analysis of Systems: 7th International Conference, TACAS ’01 Proceedings, LNCS 2031, 2001, pages 174–188.

[RS-01-3] Abstract, PostScript, PDF. Gerd Behrmann, Ansgar Fehnker, Thomas S. Hune, Kim G. Larsen, Paul Pettersson, Judi Romijn, and Frits W. Vaandrager. Minimum-Cost Reachability for Priced Timed Automata. January 2001. 22 pp. Appears in Di Benedetto and Sangiovanni-Vincentelli, editors, Hybrid Systems: Computation and Control, Fourth International Workshop, HSCC ’01 Proceedings, LNCS 2034, 2001, 147– 161.

[RS-01-2] Abstract, PostScript, PDF, DVI. Rasmus Pagh and Jakob Pagter. Optimal Time-Space Trade-Offs for Non-Comparison-Based Sorting. January 2001. ii+20 pp. Appears in Eppstein, editor, The Thirteenths Annual ACM-SIAM Symposium on Discrete Algorithms, SODA ’02 Proceedings, 2002, pages 9–18.

[RS-01-1] Abstract, PostScript, PDF, DVI. Gerth Stlting Brodal, Rolf Fagerberg, Christian N. S. Pedersen, and Anna stlin. The Complexity of Constructing Evolutionary Trees Using Experiments. January 2001. 28 pp. Appears in Orejas, Spirakis and Leeuwen, editors, 28th International Colloquium on Automata, Languages, and Programming, ICALP ’01 Proceedings, LNCS 2076, 2001, pages 140–151.